# FlashI2V: Fourier-Guided Latent Shifting Prevents Conditional Image Leakage in Image-to-Video Generation

### v1

@@ 4.1 Experimental Setup
Figure 3: Method Comparison. We compare the quantitative performance of FlashI2V (1.3B) with CogVideoX1.5-5B-I2V (Yang et al., 2024) and Wan2.1-I2V-14B-480P (Wan et al., 2025). We observe that CogVideoX1.5 and Wan2.1 exhibit color inconsistency. Additionally, Wan2.1 tends to produce extremely slow-motion or even static videos. Thanks to the avoidance of conditional image leakage, FlashI2V effectively resolves these performance degradation issues.
Training Setup. In the comparisons, we train a model for 84K steps on 20M high-quality video data collected internally, following the collection and processing pipeline described in Open-Sora Plan (Lin et al., 2024a). For each video, we randomly sample 49 frames at a fixed fps of 16, with a resolution of $\text{480}\times\text{832}$. We initialize the model from the Wan2.1-T2V-1.3B (Wan et al., 2025) model. The learnable projection is implemented using two layers of Conv3D (Tran et al., 2015) and SiLU (Elfwing et al., 2018), and the Fourier embedding layer is implemented in the same way as the patch embedding, both with zero initialization. During training, the first frame of each video serves as the conditional image. The cutoff frequency percentile of the Fourier Transform is sampled from $\mathcal{U}[0.05,0.95]$. The text prompt is dropped with a probability of 0.1. We use a batch size of 256, a learning rate of 4e-5, a weight decay of 1e-2, and the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ set to 0.999, and $\epsilon$ set to 1e-15. The weights are updated using Exponential Moving Average (EMA) with a decay of 0.9999. In the ablation study, we select a 2M subset as the training set, use a learning rate of 2e-5, a batch size of 64, and 30K training steps, while keeping the other settings unchanged.
Sampling Setup. For sampling, we use the Discrete Euler Sampler with a sigma shifting strategy as in HunyuanVideo (Kong et al., 2024), a shifting coefficient of 7.0, classifier-free guidance set to 5.0, 50 sampling steps, and a cutoff frequency percentile set to 0.1.
Evaluation. In the comparisons, we use a fixed 49 frames and the default resolution, and we utilize all Vbench-I2V (Huang et al., 2024; Zheng et al., 2025) metrics except for Camera Motion as evaluation metrics. In addition, we use ChatGPT (Achiam et al., 2023) to rewrite the short prompts of Vbench-I2V in order to obtain more accurate evaluation results. For further details, please refer to the App. A. In the ablation study, we randomly select 1,000 videos from the HD subset of OpenVid-1M (Nan et al., 2024) as the validation set, calculating the chunk-wise FVD for each setting.
@@ 4.3 Ablation Study
Generalization to out-of-domain data. We compare the chunk-wise FVD variation patterns of different I2V paradigms on in-domain and out-of-domain data, as shown in Fig. 4(a). The pseudocode implementations of the different paradigms can be found in the App. C. Each generated video is divided into four temporal chunks over time with an equal interval. Apart from FlashI2V, other paradigms concatenate the full information of the conditional image with the noisy latents. The paradigms named with "Adding Noise" add a small amount of noise to the conditional image latents, similar to the implementation in CogVideoX (Yang et al., 2024). We observe that only FlashI2V exhibits the same chunk-wise FVD variation pattern on both in-domain and out-of-domain data, indicating that the generation law learned by FlashI2V on in-domain data generalizes well to out-of-domain data. In contrast, other paradigms show inconsistent chunk-wise FVD variation patterns between in-domain and out-of-domain data, suggesting the leakage caused by shortcutting conditional images. Moreover, our method achieves the lowest FVD on out-of-domain data, meaning it has the best performance across different I2V paradigms.
@@ Appendix C Pseudo-code implementation of different I2V paradigms
Repeating Concat and Adding Noise. As shown in Algorithm 2, compared to the Repeating Concat paradigm, this approach adds a small amount of noise to the conditional images. The intensity of the noise is not strong enough to disrupt most of the information in the conditional images, but it can enhance the generalization of the model. This paradigm is used in SVD (Blattmann et al., 2023).
@@
Zero-Padding Concat and Adding Noise. As shown in Algorithm 4, compared to the Zero-padding Concat paradigm, this approach adds a small amount of noise to the conditional images. This paradigm is used in CogVideoX (Yang et al., 2024).
Inpainting. As shown in Algorithm 5, this paradigm treats I2V as a temporal completion task. For the conditional image $\bm{S}$, the approach pads $\bm{S}$ with zeros along the temporal dimension to align with the video shape. After encoding with the VAE encoder, $\bm{s}$ is obtained. Additionally, a mask is generated based on $\bm{S}$ to identify frames with information, which is then downsampled and rearranged to align with the frame numbers, height, and width of latents, resulting in $\mathbf{m}$. Both $\mathbf{m}$ and $\bm{s}$ are concatenated with noisy latents along the channel dimension as input to the denoiser. This paradigm is used in Open-Sora Plan (Lin et al., 2024a) and Wan2.1 (Wan et al., 2025).
@@ Appendix E Data Collection and Processing
Our training dataset incorporates some internal data and open-source data from Panda-70M (Chen et al., 2024c) and VIDAL (Zhu et al., 2023). We also include videos from CC0-licensed websites such as Mixkit, Pexels, and Pixabay. In addition, our data collection and processing pipeline follows the Open-Sora Plan (Lin et al., 2024a).
@@ Appendix H More Visual Results
Fig. 9 presents more inference results for FlashI2V, and the video version of the results can be found on the GitHub page.

### v2

@@ 4.1 Experimental Setup
Figure 3: Method Comparison. We compare the quantitative performance of FlashI2V (1.3B) with CogVideoX1.5-5B-I2V (cogvideox) and Wan2.1-I2V-14B-480P (wan). We observe that CogVideoX1.5 and Wan2.1 exhibit color inconsistency. Additionally, Wan2.1 tends to produce extremely slow-motion or even static videos. Thanks to the avoidance of conditional image leakage, FlashI2V effectively resolves these performance degradation issues.
Training Setup. In the comparisons, we train a model for 84K steps on 20M high-quality video data collected internally, following the collection and processing pipeline described in Open-Sora Plan (open_sora_plan). For each video, we randomly sample 49 frames at a fixed fps of 16, with a resolution of $\text{480}\times\text{832}$. We initialize the model from the Wan2.1-T2V-1.3B (wan) model. The learnable projection is implemented using two layers of Conv3D (conv3d) and SiLU (silu), and the Fourier embedding layer is implemented in the same way as the patch embedding, both with zero initialization. During training, the first frame of each video serves as the conditional image. The cutoff frequency percentile of the Fourier Transform is sampled from $\mathcal{U}[0.05,0.95]$. The text prompt is dropped with a probability of 0.1. We use a batch size of 256, a learning rate of 4e-5, a weight decay of 1e-2, and the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ set to 0.999, and $\epsilon$ set to 1e-15. The weights are updated using Exponential Moving Average (EMA) with a decay of 0.9999. In the ablation study, all models involved in the comparison are initialized from Wan2.1-T2V-1.3B. We select a 2M subset as the training set, use a learning rate of 2e-5, a batch size of 64, and 30K training steps, while keeping the other settings unchanged. Sampling Setup. For sampling, we use the Discrete Euler Sampler with a sigma shifting strategy as in HunyuanVideo (hunyuanvideo), a shifting coefficient of 7.0, classifier-free guidance set to 5.0, 50 sampling steps, and a cutoff frequency percentile set to 0.1.
Evaluation. In the comparisons, we use a fixed 49 frames and the default resolution, and we utilize all Vbench-I2V (vbench; vbench2) metrics except for Camera Motion as evaluation metrics. In addition, we use ChatGPT (gpt4) to rewrite the short prompts of Vbench-I2V in order to obtain more accurate evaluation results. For further details, please refer to the App. A. In the ablation study, we randomly select 1,000 videos from the HD subset of OpenVid-1M (openvid) as the validation set, calculating the chunk-wise FVD for each setting.
@@ 4.3 Ablation Study
Generalization to out-of-domain data. We compare the chunk-wise FVD variation patterns of different I2V paradigms on in-domain and out-of-domain data, as shown in Fig. 4(a). The pseudocode implementations of the different paradigms can be found in the App. C. Each generated video is divided into four temporal chunks over time with an equal interval. Apart from FlashI2V, other paradigms concatenate the full information of the conditional image with the noisy latents. The paradigms named with "Adding Noise" add a small amount of noise to the conditional image latents, similar to the implementation in CogVideoX (cogvideox). We observe that only FlashI2V exhibits the same chunk-wise FVD variation pattern on both in-domain and out-of-domain data, indicating that the generation law learned by FlashI2V on in-domain data generalizes well to out-of-domain data. In contrast, other paradigms show inconsistent chunk-wise FVD variation patterns between in-domain and out-of-domain data, suggesting the leakage caused by shortcutting conditional images. Moreover, our method achieves the lowest FVD on out-of-domain data, meaning it has the best performance across different I2V paradigms.
@@ 5 Conclusion
FlashI2V: Fourier-Guided Latent Shifting Prevents Conditional Image Leakage in Image-to-Video Generation
@@ Appendix C Pseudo-code implementation of different I2V paradigms
Repeating Concat and Adding Noise. As shown in Algorithm 2, compared to the Repeating Concat paradigm, this approach adds a small amount of noise to the conditional images. The intensity of the noise is not strong enough to disrupt most of the information in the conditional images, but it can enhance the generalization of the model. This paradigm is used in SVD (stable_video_diffusion).
@@
Zero-Padding Concat and Adding Noise. As shown in Algorithm 4, compared to the Zero-padding Concat paradigm, this approach adds a small amount of noise to the conditional images. This paradigm is used in CogVideoX (cogvideox).
Inpainting. As shown in Algorithm 5, this paradigm treats I2V as a temporal completion task. For the conditional image $\bm{S}$, the approach pads $\bm{S}$ with zeros along the temporal dimension to align with the video shape. After encoding with the VAE encoder, $\bm{s}$ is obtained. Additionally, a mask is generated based on $\bm{S}$ to identify frames with information, which is then downsampled and rearranged to align with the frame numbers, height, and width of latents, resulting in $\mathbf{m}$. Both $\mathbf{m}$ and $\bm{s}$ are concatenated with noisy latents along the channel dimension as input to the denoiser. This paradigm is used in Open-Sora Plan (open_sora_plan) and Wan2.1 (wan).
@@ Appendix E Data Collection and Processing
Our training dataset incorporates some internal data and open-source data from Panda-70M (panda70m) and VIDAL (languagebind). We also include videos from CC0-licensed websites such as Mixkit, Pexels, and Pixabay. In addition, our data collection and processing pipeline follows the Open-Sora Plan (open_sora_plan).
@@ Appendix H More Visual Results
Fig. 9 presents more inference results for FlashI2V, and the video version of the results can be found on the project page.